# Rapid and Safe Trajectory Planning over Diverse Scenes through Diffusion Composition

### v1

@@ III-B Diffusion Models Composition
where $N_{\text{all}}$ is the number of models need to be composed and $\mathcal{O}_{i}\in\mathcal{O}_{\text{all}}$. The obstacles $\mathcal{O}_{i}$ are encoded using a Transformer architecture as adapted from [10]. The probability density of an implicit collision-free trajectory is referred to as the unconditional probability density $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$. This distribution is directly learned from the dataset without explicit obstacle constraints. Each $p^{\text{traj}}_{\theta}$ corresponds to a diffusion model. And the $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$ and conditional diffusion model $p^{\text{traj}}_{\theta}(\bm{\tau}_{1:L}^{t}\mid\mathcal{O}_{i})$ can be achieved by adding the mask using $\mathbf{K}=[1,0]^{\top}\otimes\mathcal{O}_{i}$ to the obstacle encoder input during inference, allowing the same model to output both conditional and unconditional results.
The training of the diffusion model can be viewed as the gradient of an energy model, where holds the $\nabla_{\bm{\tau}^{t}_{1:L}}\log p^{\text{traj}}_{\theta}(\bm{\tau}^{t}_{1:L}\mid\mathcal{O}_{i})\propto\epsilon^{\text{traj}}_{\theta}(\bm{\tau}^{t}_{1:L},t,\mathcal{O}_{i})$. Thus, Eq. (2) can be rewritten as:
@@
III-C Safety Filter
In test-time planning, denoising from random Gaussian noise introduces variations in the planned trajectories, which can potentially cause instability and even collisions. To ensure the safety of the planned trajectory, after a batch of trajectories is generated, the one that satisfies both strong control feasibility and safety is selected as the final trajectory based on specific metrics. The overall process is in Algorithm 1.
The batch of all trajectories is denoted as $\mathcal{B}^{\text{test}}=\{\mathbf{q}_{i},\bm{\varphi}_{i}\}_{i=1}^{N_{\text{filter}}}$, where $\mathbf{q}_{i}=[x,\,y]^{\top}_{2\times L}$ represents the points of the $i$-th trajectory, and $N_{\text{filter}}$ is the number of trajectories. The $\mathcal{B}^{\text{test}}_{i}=\{\mathbf{q}_{i},\bm{\varphi}_{i}\}$ can be derivatived from the planned trajectory $\bm{\tau}_{i}=[x_{i},y_{i},q_{i,z},q_{i,w}]^{\top}_{4\times L}$.
First, the control feasibility is considered. The distance between adjacent points in $\mathbf{q}_{i}$ is given by: $\mathbf{d}_{i}={\left(\Delta\mathbf{q}_{i,[1,:]}^{2}+\Delta\mathbf{q}_{i,[2,:]}^{2}\right)^{\frac{1}{2}}}$, where $\Delta\mathbf{q}_{i}=\mathbf{q}_{i,[:,2:L]}-\mathbf{q}_{i,[:,1:L-1]}$. Then the trajectory length of $\mathbf{q}_{i}$ can be given by $l_{i}=\sum_{j=1}^{L-1}d_{i,j}$. It is used as a metric because minimizing the trajectory length helps to avoid redundant GSPs. The velocity $\mathbf{v}_{i}$ and acceleration $\mathbf{a}_{i}$ are then calculated as: $\mathbf{v}_{i}=\frac{\mathbf{d}_{i}}{\Delta t}\,\text{and}\,\mathbf{a}_{i}={(\mathbf{v}_{i+1}-\mathbf{v}_{i})}/{\Delta t}$. Then the yaw rate is given by: $\mathbf{r}={\left(\bm{\varphi}_{i,[2:L]}-\bm{\varphi}_{i,[1:L-1]}\right)}/{\Delta t}$. Finally, the corresponding steering angle is: $\bm{\delta}_{i}=\arctan\left(\frac{l_{w}\cdot\mathbf{r}_{i}}{\mathbf{v}_{i}}\right)$, where $l_{w}$ is the wheelbase of the vehicle.
To guarantee numerical stability, we normalize the evaluation terms so their scales remain uniform. After obtaining $\mathbf{L}=\{l_{i}\}_{i=1}^{N_{\text{filter}}}$, $\mathbf{a}=\{\mathbf{a}_{i}\}_{i=1}^{N_{\text{filter}}}$, and $\bm{\delta}=\{\bm{\delta}_{i}\}_{i=1}^{N_{\text{filter}}}$ for the trajectory batch $\mathcal{B}^{\text{test}}$, the corresponding values are normalized. The normalization process for $\mathbf{L}$ is given by $\frac{\mathbf{L}-\min(\mathbf{L})}{\max(\mathbf{L})-\min(\mathbf{L})}.$ For acceleration batch $\mathbf{a}$, the normalization is: $\frac{\mathbf{\|a\|}^{2}-\min(\mathbf{\|a\|}^{2})}{\max(\mathbf{\|a\|}^{2})-\min(\mathbf{\|a\|}^{2})}.$ The same normalization process is applied to $\bm{\delta}$. Then the final normalized values are denoted as $\tilde{\mathbf{L}}$, $\tilde{\mathbf{a}}$, and $\tilde{\bm{\delta}}$, respectively.
@@ III-C Safety Filter
$J_{i}=\tilde{\mathbf{L}}_{i}+\|\tilde{\mathbf{a}}_{i}\|+\|\tilde{\bm{\delta}}_{i}\|+J_{i,\text{safe}},\forall\,i\in[1,N_{\text{filter}}]$ (4)
where $J_{\text{best}}=\min\{J_{i}\}_{i=1}^{N_{\text{filter}}}$. The smaller the value of $J_{i}$, the higher the quality of the trajectory. The final computed value of $J_{\text{best}}$ corresponds to a single trajectory $\mathcal{B}$, which exhibits both safety and strong control feasibility.
To prevent abrupt changes in the yaw angle along the selected trajectory $\mathcal{B}$, a posterior check is conducted on the variation $\Delta\bm{\varphi}=\bm{\varphi}_{[2:L]}-\bm{\varphi}_{[1:L-1]}$ between adjacent trajectory points. If any $\Delta\bm{\varphi}$ exceeds $w_{max}\cdot\Delta t$, where $w_{\max}$ denotes the vehicle’s maximum angular velocity, the corresponding metric $\phi(\mathcal{B})$ is assigned a significant penalty value $V_{\text{inf}}$; otherwise, it is set to 0. This mechanism is designed to filter out trajectories that violate motion constraints, thereby ensuring control feasibility, which is denoted as follows:
@@
This safety filter simultaneously accounts for both control feasibility and safety, thereby enhancing the overall reliability and robustness of the planned trajectories.
@@
7:        Return; % Find trajectory with safety and control
8:         feasibility.
9:     end if
10:     $k\leftarrow k+1$; % Replanning.
11:  end while
@@
IV EXPERIMENTAL RESULTS AND DISCUSSION
The proposed RSTP method is validated through both simulation and real-world tests. In simulation, individual model and their compositions are evaluated using safety, computational efficiency, and control feasibility metrics. The safety filter’s effectiveness is also confirmed in simulation. Trajectories generated by the composed models are then tested on the F1TENTH vehicle platform and show strong safety and control feasibility under real-world uncertainties.
IV-A Environments and Baselines
The following scenes and experimental setups are established for simulation and real-world testing:
@@ IV-A Environments and Baselines
    Static Scene ($SS$): Includes only static obstacles. The diffusion model trained on this scene is referred to as the Static Model (StM).
@@
[…]
@@
[…]
The architecture and hyperparameters of the diffusion models follow the open-source implementation [^1] https://github.com/devinluo27/potential-motion-plan-release, with the composition formula aligned to (3). The more detailed setup is in Appendix -A.
@@
Methos4 F.Rate1 C.Rate2 C.T M.TE Max Min Mean M.RP Std Max Min Mean M.RP Std Proposed RSTP 0.57% 8.80% 0.522 0.137 0.210 98.88% 3.16e-4 12.909 0.321 1.424 47.23% 1.647 RSTP-NSF 4.50% 11.87% 0.515 0.144 0.195 98.96% 1.65e-4 37.933 0.284 1.495 44.62% 4.299 Heuristics $\text{Hybrid A}^{\star}$ 24.88% 27.75% 248.964 0.004 18.747 - $\uparrow$ 6.929 0.459 2.699 - 0.882 Kino RRT 14.64% 58.18% 39.362 0.019 4.267 77.24% 31.601 39.686 0.476 3.232 -19.77% 16.876 $\text{A}^{\star}$ 0.86% 67.37% 5.337 0.138 2.032 89.16% 0.924 53.129 1.085 3.903 -44.62% 8.772 $\text{RRT}^{\star}$ 1.53% 59.43% 13.950 0.160 3.177 83.05% 3.029 23.561 0.951 4.290 -58.96% 12.235 $\text{FMT}^{\star}$ 0.77% 74.07% 2.766 0.615 1.450 92.27% 0.124 10.194 0.284 5.188 -92.23% 4.237 Optimization3 RITP 0.76% 7.87% 0.351 0.008 0.059 99.69% 0.002 6.413 0.254 0.792 70.67% 0.377 OBCA 10.41% 13.20% 1.395 0.378 0.508 97.29% 0.005 5.877 0.178 0.560 79.27% 0.111
@@
[…]
IV-B Evaluate the Performance of RSTP Method in Simulation
@@
IV-B1 On the Safety Performance of our RSTP method
@@
The results summarized in Table I show that the proposed RSTP achieves a Failure Rate (F.Rate) of only 0.57% and a low Collision Rate (C.Rate) of 8.8% during tracking. By implicitly learning the vehicle kinematics constraints, RSTP significantly reduces collisions during tracking, outperforming kinematics-aware methods such as Hybrid $\text{A}^{\star}$ and Kino RRT, which suffer from higher failure rates of 24.88% and 14.64% due to the limitations of discrete search and sampling in long-horizon planning. Optimization-based methods, RITP and OBCA, explicitly enforce kinematics and achieve comparable collision rates of 7.87% and 13.2%, compared to RSTP. However, they require well-defined warm-start paths, which increase computation time. More analysis on the control feasibility of planned trajectories is available at Appendix -B. In summary, RSTP delivers notable safety and efficiency gains, benefits to the diffusion model’s generative ability, and stability.
IV-B2 On the time efficiency of our RSTP method
@@
As shown in Table I, the proposed RSTP achieves a mean C.T of only 0.21s of all starting poses in the evaluation dataset $D_{A}$, representing a 98.88% improvement of computational efficiency compared to BLM. Additionally, it demonstrates strong stability, with a standard deviation of only 3.16e-4, outperforming all comparison methods. This time efficiency stems from state-based denoising, which significantly reduces the noise dimension. The time stability benefits from the strong stability of diffusion models in generating trajectories. Optimization-based methods, such as RITP and OBCA, depend on Hybrid $\text{A}^{\star}$ for initialization. This dependency results in an additional mean computational cost of 18.747 s. In contrast, the RSTP method avoids this extra warm-start overhead. In conclusion, state-based diffusion inference facilitates rapid trajectory planning, enabling the vehicle to respond quickly to environmental conditions with strong stability.
IV-B3 On the effectiveness of Safety Filter

### v2

@@ IV-A Diffusion Models Composition
where $N_{\text{all}}$ is the number of models to be composed and $\mathcal{O}_{i}\in\mathcal{O}_{\text{all}}$. The obstacles $\mathcal{O}_{i}$ are encoded using a Transformer architecture as adapted from [10]. The probability density of a collision-free trajectory is referred to as the unconditional probability density $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$. This distribution is directly learned from the dataset without explicit obstacle constraints. Each $p^{\text{traj}}_{\theta}$ corresponds to a diffusion model. And the $p^{\text{traj}}_{\theta_{\text{uncond}}}(\bm{\tau}_{1:L}^{t})$ and conditional diffusion model $p^{\text{traj}}_{\theta}(\bm{\tau}_{1:L}^{t}\mid\mathcal{O}_{i})$ can be obtained by masking the obstacle-encoder input with $\mathbf{K}=[1,0]^{\top}\otimes\mathcal{O}_{i}$ during inference, allowing the same model to output both conditional and unconditional results.
The training of the diffusion model can be viewed as the gradient of an energy model, where satisfies $\nabla_{\bm{\tau}^{t}_{1:L}}\log p^{\text{traj}}_{\theta}(\bm{\tau}^{t}_{1:L}\mid\mathcal{O}_{i})\propto\epsilon^{\text{traj}}_{\theta}(\bm{\tau}^{t}_{1:L},t,\mathcal{O}_{i})$. Thus, Eq. (2) can be rewritten as:
@@
IV-B Safety Filter
In test-time planning, denoising from random Gaussian noise introduces variations in the planned trajectories, which can potentially cause instability and even collisions. To ensure safety, a batch of trajectories is generated, and the highest-quality one is selected as the final trajectory. The overall process is in Algorithm 1.
The batch of all trajectories is denoted as $\mathcal{B}^{\text{test}}=\{\mathbf{q}_{i},\bm{\varphi}_{i}\}_{i=1}^{N_{\text{filter}}}$, where $\mathbf{q}_{i}=[x,\,y]^{\top}_{2\times L}$ denotes the point sequence of the $i$-th trajectory, and $N_{\text{filter}}$ is the number of trajectories. The $\mathcal{B}^{\text{test}}_{i}=\{\mathbf{q}_{i},\bm{\varphi}_{i}\}$ can be derived from the planned trajectory $\bm{\tau}_{i}=[x_{i},y_{i},q_{i,z},q_{i,w}]^{\top}_{4\times L}$.
First, the kinematic feasibility is considered. The distance between adjacent points in $\mathbf{q}_{i}$ is given by: $\mathbf{d}_{i}={\left(\Delta\mathbf{q}_{i,[1,:]}^{2}+\Delta\mathbf{q}_{i,[2,:]}^{2}\right)^{\frac{1}{2}}}$, where $\Delta\mathbf{q}_{i}=\mathbf{q}_{i,[:,2:L]}-\mathbf{q}_{i,[:,1:L-1]}$. Then the trajectory length of $\mathbf{q}_{i}$ can be given by $l_{i}=\sum_{j=1}^{L-1}d_{i,j}$. It is used as a metric because minimizing the trajectory length helps to avoid redundant GSPs. The velocity $\mathbf{v}_{i}$ and acceleration $\mathbf{a}_{i}$ are then calculated as: $\mathbf{v}_{i}=\frac{\mathbf{d}_{i}}{\Delta t}\,\text{and}\,\mathbf{a}_{i}={(\mathbf{v}_{i+1}-\mathbf{v}_{i})}/{\Delta t}$. Then the yaw rate is given by: $\mathbf{r}={\left(\bm{\varphi}_{i,[2:L]}-\bm{\varphi}_{i,[1:L-1]}\right)}/{\Delta t}$. Finally, the corresponding steering angle is: $\bm{\delta}_{i}=\arctan\left(\frac{l_{w}\cdot\mathbf{r}_{i}}{\mathbf{v}_{i}}\right)$, where $l_{w}$ is the wheelbase of the vehicle.
To guarantee numerical stability, we normalize the cost terms so their scales remain uniform. After obtaining $\mathbf{L}=\{l_{i}\}_{i=1}^{N_{\text{filter}}}$, $\mathbf{a}=\{\mathbf{a}_{i}\}_{i=1}^{N_{\text{filter}}}$, and $\bm{\delta}=\{\bm{\delta}_{i}\}_{i=1}^{N_{\text{filter}}}$ for the trajectory batch $\mathcal{B}^{\text{test}}$, the corresponding values are normalized. The normalization process for $\mathbf{L}$ is given by $\frac{\mathbf{L}-\min(\mathbf{L})}{\max(\mathbf{L})-\min(\mathbf{L})}.$ For acceleration batch $\mathbf{a}$, the normalization is: $\frac{\mathbf{\|a\|}^{2}-\min(\mathbf{\|a\|}^{2})}{\max(\mathbf{\|a\|}^{2})-\min(\mathbf{\|a\|}^{2})}.$ The same normalization process is applied to $\bm{\delta}$. Then the final normalized values are denoted as $\tilde{\mathbf{L}}$, $\tilde{\mathbf{a}}$, and $\tilde{\bm{\delta}}$, respectively.
@@ IV-B Safety Filter
$J_{i}=\omega_{1}\tilde{\mathbf{L}}_{i}+\omega_{2}\|\tilde{\mathbf{a}}_{i}\|+\omega_{3}\|\tilde{\bm{\delta}}_{i}\|+\omega_{4}J_{i,\text{safe}},\forall\,i\in[1,N_{\text{filter}}]$ (4)
where $\omega_{1}$ to $\omega_{4}$ are the corresponding weights. Each of the cost terms in (4) is normalized, allowing their corresponding weights to be easily tuned in different scenarios. $J_{\text{best}}=\min\{J_{i}\}_{i=1}^{N_{\text{filter}}}$. The smaller the value of $J_{i}$, the higher the quality of the trajectory. The final computed value of $J_{\text{best}}$ corresponds to a single trajectory $\mathcal{B}$, which demonstrates superior safety, smoothness, and efficiency among a batch of candidate trajectories.
To prevent abrupt changes in the yaw angle along the selected trajectory $\mathcal{B}$, a post hoc check is conducted on the variation $\Delta\bm{\varphi}=\bm{\varphi}_{[2:L]}-\bm{\varphi}_{[1:L-1]}$ between adjacent trajectory points. If any $\Delta\bm{\varphi}$ exceeds $w_{max}\cdot\Delta t$, where $w_{\max}$ denotes the vehicle’s maximum angular velocity, the corresponding metric $\phi(\mathcal{B})$ is assigned a significant penalty value $V_{\text{inf}}$; otherwise, it is set to 0. This mechanism is designed to filter out trajectories that violate motion constraints, thereby ensuring kinematic feasibility, which is denoted as follows:
@@
For execution safety, if the number of replanning attempts exceeds $N_{\text{retry}}$ without finding a feasible solution, the planner returns an empty result and the vehicle refrains from acting.
@@
7:   Return $\tau^{\star}$; % Find the optimal trajectory.
8:  end if
9:  $k\leftarrow k+1$; % Replanning.
10:end while
11: Return $\emptyset$;
@@
V EXPERIMENTAL RESULTS AND DISCUSSION
The proposed RSTP method is validated through both simulation and real-world tests. In simulation, individual model and their compositions are evaluated using safety, computational efficiency, and kinematic feasibility metrics. The safety filter’s effectiveness is also confirmed in simulation. Trajectories generated by the composed models are then tested on the F1TENTH vehicle platform and show strong safety and kinematic feasibility under real-world uncertainties.
V-A Environments and Baselines
Details of the experimental setup are provided in Appendix A. The diffusion model adopts the architecture and hyperparameters of [10], and its composition formula follows (3). In all experiments, the weights of the terms in Eq.(4) are set equal. The following scenes and experimental setups are established for simulation and real-world testing:
@@ V-A Environments and Baselines
    Static Scene ($SS$): Includes only static obstacles. The diffusion model trained on the static scene is referred to as the Static Model (StM).
@@
    Composed Scene 1 ($CS_{1}$): Features the same static obstacles as SS, along with a single moving obstacle vehicle. This vehicle is initialized at $x_{1}=3.5,y_{1}=1.2,q_{z}=1.0,q_{w}=0.00]$ and moves in a straight line at a constant speed of $0.4\text{m/s}$ for $7.1\,\text{s}$, as shown in Fig. 6(a). The diffusion model trained only on this dynamic obstacle trajectory in $CS_{1}$ is denoted as Dynamic Model 1 (DyM1).
@@
    Composed Scene 2 ($CS_{2}$): Features the same static obstacles as SS, along with a single moving obstacle vehicle. This vehicle is initialized at $x_{2}=3.5,y_{2}=2.0,q_{z}=0.891,q_{w}=-0.454]$ and moves in a straight line at a constant speed of $0.4\text{m/s}$ for $5.3\,\text{s}$, as shown in Fig. 6(b). The diffusion model trained only on this dynamic obstacle trajectory in $CS_{2}$ is denoted as Dynamic Model 2 (DyM2).
We compare our RSTP method with multiple existing approaches. Baseline methods for comparison include $\text{A}^{\star}$[14], $\text{RRT}^{\star}$[15], and $\text{FMT}^{\star}$[16], all of which disregard vehicle kinematics. Additionally, kinematics-aware methods, Kino RRT[17] and $\text{Hybrid A}^{\star}$[1] are included for comparison. We denote the baseline method $\text{Hybrid A}^{\star}$ as BLM. The optimization-based methods evaluated are Rapid Iterative Trajectory Planning (RITP) [3] and Optimization-Based Collision Avoidance (OBCA) [2], both of which are initialized using paths generated by $\text{Hybrid A}^{\star}$. To assess kinematic feasibility, the MPC tracker described in Section IV is employed as the tracking controller. The mean tracking error (M.TE) serves as the metric for evaluating kinematic feasibility. Computational efficiency is evaluated by measuring the computation time (C.T). For the diffusion model, the C.T is the time to generate the final collision-free trajectory, including replanning. The percentage reduction in C.T compared to the BLM is denoted as $\text{M.RP}=\frac{\text{C.T}_{\text{BLM}}-\text{C.T}}{\text{C.T}_{\text{BLM}}}$. Similarly, the percentage reduction in M.TE is represented by $\text{M.RP}=\frac{\text{M.TE}_{\text{BLM}}-\text{M.TE}}{\text{M.TE}_{\text{BLM}}}$. The configuration with the safety filter disabled is indicated as RSTP-NSF.
@@
Methods4 F.Rate1 C.Rate2 C.T M.TE Max Min Mean M.RP Std Max Min Mean M.RP Std Proposed RSTP 0.57% 8.80% 0.522 0.137 0.210 98.88% 3.16e-4 12.909 0.321 1.424 47.23% 1.647 RSTP-NSF 4.50% 11.87% 0.515 0.144 0.195 98.96% 1.65e-4 37.933 0.284 1.495 44.62% 4.299 Heuristics $\text{Hybrid A}^{\star}$ 24.88% 27.75% 248.964 0.004 18.747 - $\uparrow$ 6.929 0.459 2.699 - 0.882 Kino RRT 14.64% 58.18% 39.362 0.019 4.267 77.24% 31.601 39.686 0.476 3.232 -19.77% 16.876 $\text{A}^{\star}$ 0.86% 67.37% 5.337 0.138 2.032 89.16% 0.924 53.129 1.085 3.903 -44.62% 8.772 $\text{RRT}^{\star}$ 1.53% 59.43% 13.950 0.160 3.177 83.05% 3.029 23.561 0.951 4.290 -58.96% 12.235 $\text{FMT}^{\star}$ 0.77% 74.07% 2.766 0.615 1.450 92.27% 0.124 10.194 0.284 5.188 -92.23% 4.237 Optimization3 RITP 0.76% 7.87% 0.351 0.008 0.059 99.69% 0.002 6.413 0.254 0.792 70.67% 0.377 OBCA 10.41% 13.20% 1.395 0.378 0.508 97.29% 0.005 5.877 0.178 0.560 79.27% 0.111
@@
V-B Evaluate the Performance of RSTP Method in Simulation
@@
V-B1 On the Safety Performance of our RSTP method
@@
The results summarized in Table I show that the proposed RSTP achieves a Failure Rate (F.Rate) of only 0.57% and a low Collision Rate (C.Rate) of 8.8% during tracking. By implicitly learning the vehicle kinematics constraints, RSTP significantly reduces collisions during tracking, outperforming kinematics-aware methods such as Hybrid $\text{A}^{\star}$ and Kino RRT, which suffer from higher failure rates of 24.88% and 14.64% due to the limitations of discrete search and sampling in long-horizon planning. Optimization-based methods, RITP and OBCA, explicitly enforce kinematics and achieve comparable collision rates of 7.87% and 13.2%, compared to RSTP. However, they require well-defined warm-start paths, which increase computation time. More analysis on the kinematic feasibility of planned trajectories is available at Appendix -B. In summary, RSTP delivers notable safety and efficiency gains, benefits to the diffusion model’s generative ability, and stability.
V-B2 On the Time Efficiency of our RSTP method
@@
As shown in Table I, the proposed RSTP achieves a mean C.T of only 0.21s across all starting poses in the evaluation dataset $D_{A}$, representing a 98.88% improvement of computational efficiency compared to BLM. Additionally, it demonstrates strong stability, with a standard deviation of only 3.16e-4, outperforming all comparison methods. This time efficiency stems from state-based denoising, which significantly reduces the noise dimension. The time stability benefits from the strong stability of diffusion models in generating trajectories. Optimization-based methods, such as RITP and OBCA, depend on Hybrid $\text{A}^{\star}$ for initialization. This dependency results in an additional mean computational cost of 18.747 s. In contrast, the RSTP method avoids this extra warm-start overhead. In conclusion, state-based diffusion inference facilitates rapid trajectory planning, enabling the vehicle to respond quickly to environmental conditions with strong stability.
V-B3 On the Effectiveness of Safety Filter